OpenStudy (anonymous):

Please, Please, Please Help Me! How do I calculate the wavelength (λ) of transitions between the n-orbital? For example: \(\ \Huge n=7 \rightarrow n=2 .\)

OpenStudy (anonymous):

\[\frac{ 1 }{ \lambda } = R(\frac{ 1 }{ S^{2} }-\frac{ 1 }{ L ^{2} })\] Where R is Rydberg's constant S is energy level transitions to L is energy level transitioned from

OpenStudy (anonymous):

? What do I do with 1/λ???

OpenStudy (anonymous):

Once you've plugged in your values and found what 1/λ equals just take 1/(that value) which will give you λ